# Beyond Browsing: API-Based Web Agents

### v1

@@ 5.2 API Statistics for WebArena Sites
In this section, we provide a detailed analysis of the API support for various websites used in the WebArena tasks, categorized into three levels: good, medium, and poor. The availability, functionality, and documentation of APIs, as described in Table 1, play a crucial role in the efficiency and flexibility of our agents.
| Websites | Gitlab | Map | Shopping | Admin | |
| Number of Endpoints | 988 | 53 | 556 | 556 | 31 |
| API/Doc Quality | Good | Good | Fair | Fair | Poor |
@@ 5.2.1 Good API Support
For Gitlab, we leveraged the open Gitlab REST APIs111111Documentation of all Gitlab APIs could be found at https://docs.gitlab.com/ee/api/rest/., which consist of 988 endpoints. These APIs offer extensive coverage across a wide range of functionalities, including repositories, commits, users, merge requests, and issues. This comprehensive API support allows for effective interaction with most tasks required in WebArena, making it one of the best-supported platforms in terms of API availability.
The majority of Gitlab-related tasks can be handled with the provided APIs, with only a small fraction of tasks, such as retrieving the user’s Gitlab feed token, not covered by any existing endpoints. Overall, Gitlab’s API structure provides robust support.
Map
The Map website offers three sets of APIs, each offering distinct functionalities, with a total of 53 endpoints. Although fewer in number compared to Gitlab and Shopping, these APIs still provide significant coverage for the tasks in WebArena.
The first set of APIs, openly available at Nominatim121212https://nominatim.org/release-docs/develop/api/Overview/, offers essential endpoints for geographic searches. The second set of APIs, from Project OSRM131313Openly available at https://project-osrm.org/docs/v5.5.1/api, focuses on routing and navigation functionalities. The third set of APIs, available at OpenStreetMap141414Publicly available at https://wiki.openstreetmap.org/wiki/API_v0.6, deals primarily with map database operations. This API is rarely used in WebArena tasks but offers capabilities for interacting with OSM data.
Despite the smaller number of endpoints compared to other websites, the APIs available for the Map tasks are mostly well-documented and cover most of the essential WebArena use cases.
@@ 5.2.2 Medium API Support
The Shopping and Shopping Admin websites share a common set of APIs from the Adobe Commerce API151515https://developer.adobe.com/commerce/webapi/rest/quick-reference/, consisting of 556 endpoints. These APIs provide a reasonable level of support for common shopping tasks such as managing products, categories, and customer accounts.
However, some features are absent, such as the ability to add items to a wish list, and thus these tasks must be handled via web browsing. Despite this, the API documentation is fairly detailed and covers most core functionalities, making it a solid, though not exhaustive, solution for handling shopping-related tasks.
@@ 5.2.3 Poor API Support
The Reddit tasks in WebArena are based on a self-hosted limited clone of the Reddit website 161616https://codeberg.org/Postmill/Postmill, with limited functionalities as compared to the official site. As a result, all of the available APIs are self-implemented, with a best effort to mimic to official Reddit APIs. With only 31 endpoints, this website offers minimal API support and no API documentation, making it the least API-friendly website in the benchmark.
Many critical functionalities, such as searching for specific posts, are missing, leaving agents to rely heavily on web browsing to complete tasks. The limited API support significantly hampers the efficiency of task execution on Reddit, highlighting the need for a hybrid browsing+API approach.
@@ 5.3 API Implementation Details
In this section, we will discuss how we provided the APIs to the agents when evaluating different web applications inside WebArena, where we follow the methodologies as discussed in Section 3.3.
@@ 5.3.1 One-Stage Documentation for Small API Sets
For websites with fewer than 100 API endpoints, namely the Map and Reddit websites, we directly incorporated the full documentation into the prompt provided to the agent.
In the case of the Map API, the documentation was sourced directly from the public API documentation provided for the website. The only modification made was the addition of an explanation detailing how to make HTTP requests using the requests library in Python for interacting with the Map API’s endpoints. This ensured that the agent could comprehend both the structure of the API and how to implement calls programmatically.
For Reddit, since there was no pre-existing documentation for the APIs, we leveraged GPT-4o171717https://openai.com/index/hello-gpt-4o/ itself to generate these README files. By prompting GPT-4o with a file containing all implementations of the API endpoints, we generated a README documentation, including input parameters, expected outputs, and example API calls.
@@ 5.3.2 Two-Stage Documentation Retrieval for Large API Sets
For websites with more than 100 endpoints, such as GitLab, Shopping, and Shopping Admin, we employ a two-stage documentation retrieval process.
For GitLab, we obtained the README documentation from the official GitLab REST API documentation site. For the Shopping and Shopping Admin websites, the documentation was provided in the form of an OpenAPI specification, structured in YAML format.
@@ 5.4 Evaluation Framework
We employed OpenHands as our primary evaluation framework to facilitate the development and testing of our agents (Wang et al., 2024c). OpenHands is an open-source platform designed for creating and evaluating AI agents that interact with both software and web environments, making it an appropriate infrastructure for our proposed methods. The OpenHands architecture supports a variety of interfaces for agents to interact with. Moreover, this framework allows agents to keep a detailed record of past actions in the prompt, enabling agents to execute actions in a way that is consistent with earlier steps. For coding tasks, it implements an agent based on CodeAct (Wang et al., 2024a) that incorporates a sandboxed bash operating system and Jupyter IPython181818https://ipython.org environments, enabling the execution of Python code. Additionally, it includes a BrowsingAgent browsing agent that focuses solely on web navigation. This agent operates within a Chromium web browser powered by Playwright191919https://playwright.dev/, utilizing a comprehensive set of browser actions defined by BrowserGym (Drouin et al., 2024b). However, while the browsing agent can browse websites, and the CodeActAgent make API calls and execute code, there is not an agent that can natively do both. Given this base, we developed two varieties of agents for API-based solving of web tasks.
API-Based Agent
First, our API-based agent essentially uses the CodeAct architecture (Wang et al., 2024a). In addition to the basic CodeAct framework, we tailor the agent for API calling by adding specialized instructions and examples that guide its understanding of various API endpoints and their usage. At each step, the agent could utilize all previous actions to make informed selection of actions. The prompt of the API-Based Agent is included in the Appendix A.2.
Hybrid Browsing/API Calling Agent
In addition to the API-based agent, we developed a Hybrid Agent that integrates Chromium web browsing functionalities powered by Playwright into the existing framework of the API-based agent. This Hybrid Agent is provided the prompt describing both the APIs and the browsing actions, allowing for free transitions between API calling and web browsing. At each step, the agent can utilize the current state of the browser, all previous actions taken by the agent, and the results of those actions to determine the next course of action. The prompt of the Hybrid Agent is included in the Appendix A.3.
6 Results
6.1 Main Results
@@
| WebArena Base (Zhou et al., 2023) | 15.0 | 15.6 | 13.9 | 10.4 | 6.6 | 8.3 | 12.3 |
@@
| Hybrid Agent | 44.4 | 45.9 | 25.7 | 41.2 | 28.3 | 16.7 | 35.8 |
@@
The main results of our evaluation, as summarized in Table 2, demonstrate the performance of three different agents across the websites in the WebArena benchmark.
@@
The API-Based Agent consistently performed well across most tasks, achieving higher scores in all websites compared to the Browsing agent. This agent’s strong performance is attributed to its specialized design for API calling, enabling it to efficiently interact with the websites’ APIs and complete tasks with minimal reliance on browsing capabilities.
@@
In contrast, the Browsing Agent, which is designed solely for navigating web interfaces, demonstrated significantly lower performance across all domains. It achieved its best scores on Gitlab and Map, but struggled more on Reddit.
@@
The Hybrid Agent, which integrates both API calling and web browsing, outperformed the other agents in all categories. The agent’s ability to dynamically switch between API calling and web browsing proved beneficial. API calling delivers high performance for web tasks when well-supported APIs are available, while web browsing serves as a backup when API endpoints are unavailable or incomplete. Even if the website provides comprehensive APIs, there might be corner cases where APIs are not supportive. In these cases, relying on web browsing is still needed for tasks that would otherwise fail through API-only interactions. Table 3 documents the percentage of actions of our Hybrid Agent. Across all websites, our Hybrid Agent chooses to do both Browsing and API in the same task at least half of the time.
Table 4 documents the accuracy of the Hybrid Agent across website when performing different choices of actions. We can see that it show consistently high accuracy when choosing API only and API+browsing.
| Actions | Gitlab | Map | Shopping | Admin | Multi | AVG. | |
| Browsing only | 7.8 | 3.7 | 38.5 | 2.2 | 17.0 | 8.3 | 14.3 |
| API only | 21.1 | 4.6 | 7.5 | 1.1 | 0.9 | 10.4 | 8.0 |
| Browsing+API | 71.1 | 91.7 | 54.0 | 96.7 | 82.1 | 81.3 | 77.7 |
| Choices of Action | Gitlab | Map | Shopping | Admin | Multi | AVG. | |
| Browsing only | 7.1(1/14) | 50.0(2/4) | 23.6(17/72) | 50.0(2/4) | 11.1(2/18) | 25.0(1/4) | 21.6(25/116) |
| API only | 47.4(18/38) | 40.0(2/5) | 21.4(3/14) | 50.0(1/2) | 0.0(0/1) | 20.0(1/5) | 38.5(25/65) |
| Browsing+API | 47.7(61/128) | 46.0(46/100) | 27.7(28/101) | 40.9(72/176) | 32.2(28/87) | 15.4(6/39) | 38.2(241/631) |
@@
| Browsing Agent | API-Based Agent | Hybrid Agent | |||
| steps | cost | steps | cost | steps | cost |
| 8.4 | $0.1 | 7.8 | $1.2 | 8.9 | $1.5 |
Additionally, we use Table 5 to demonstrate the average steps taken and the average cost for each agent to complete WebArena tasks. The breakdown of steps and cost by website is in the Appendix A.4. Figure 4 demonstrates a scatterplot of the average accuracy of each agent on WebArena over their average steps and average cost.
Steps
The browsing agent consistently takes more steps to complete tasks compared to the API-based agent, while the Hybrid Agent takes the most steps amongst the three agents. This is likely due to the browsing agent’s reliance on navigating web interfaces and interacting with visual elements, which involves a sequential and more time consuming processes. The API-based agent is the most efficient in terms of steps, as it can directly interact with structured data via APIs, bypassing many of the steps involved in traditional web navigation. The Hybrid Agent, combining both action spaces from the browsing agent and the API-based agent, takes more steps than both agents.
Costs
The cost of completing tasks shows a different trend. While the browsing agent requires more steps, it is much cheaper compared to the API-based agent and the Hybrid Agent. This is primarily because the prompts needed for browsing agents are much shorter. When browsing, the agent only needs instructions on how to use the web interface and the limited action space around 14 browsing actions. In contrast, API-based and Hybrid Agents require access to a much larger set of API calls. For example, when interacting with GitLab, the agent is provided with 988 available APIs, leading to much longer prompts and significantly increasing the cost of execution. The cost goes down when the prompt for API calling is shorter. For example, the Reddit website has the least length of API documentation, where its cost is also less than other websites. However, as visualized in Figure 4, the accuracy of the API-based agent and the Hybrid Agent is much higher than the browsing agent, which makes the increase in cost justifiable due to the significantly improved task performance. The higher cost is offset by the agents’ ability to complete tasks more accurately and efficiently. In the future, this increased cost could potentially be mitigated by methods that retrieve only relevant APIs on the fly.
@@ 6.2 Does API Quality Matter?
The short answer is yes, API quality does significantly impact the performance of the API-based agent. High quality APIs provide comprehensive and well-documented endpoints that enable agents to interact accurately and efficiently with websites. With comprehensive API support, the API-based agent is able to tackle more tasks through API calling, while the Hybrid Agent could rely less on the browsing agent; on the other hand, clear and detailed documentation allows agents to utilize the APIs effectively, ensuring that requests are accurate, and minimizing potential errors in task execution. For example, the websites Gitlab and Map with the best API support as mentioned in Section 5.2, demonstrates the highest task completion accuracy by the API-based agent and the Hybrid Agent across all websites.
Conversely, low-quality APIs, characterized by incomplete functionality or ambiguous documentation, can significantly degrade performance. In such cases, the absence of necessary endpoints may prevent the API-based agent from completing tasks, forcing the Hybrid Agent to resort to web browsing. Moreover, poorly documented APIs can result in incorrect parameters and headers being used, further reducing the effectiveness of the agent. This highlights the importance for websites to maintain comprehensive and well-documented API support.
@@
6.3 Case Studies
Case 1
One example where the Hybrid Agent succeeded, while both the API-based and browsing agents failed, involved a task from the Shopping Admin domain. The query was to ”delete all negative reviews for Sybil running short,” a product listed in the shopping admin interface. In this instance, the API-based agent failed because no relevant API endpoints were available for retrieving or deleting reviews. Similarly, the browsing agent failed, as completing this task purely through web navigation required too many steps, as depicted in Figure 5. This complexity made the task challenging for an agent relying solely on web interactions. However, the Hybrid Agent successfully completed the task by leveraging both API and browsing functionalities. An example trace of the Hybrid Agent shown in Figure 5. This case highlights the Hybrid Agent’s ability to efficiently combine API calls with web interactions, allowing it to tackle complex multi-step tasks that would be difficult or impossible for solely browsing or solely API-based agents.
Case 2
7 Related Work
The development of AI agents that interact with the web and APIs has garnered significant research attention. Web browsers, serving as the primary interface for interacting with online content, have long been a focus for AI research. Web-based agents that can navigate websites, extract information, and perform tasks autonomously have been studied extensively, especially in the context of large language models (LLMs) and agents designed to mimic human behavior online.
Web Navigation Agents
Much prior work has centered around agents that perform web-based tasks using browsing actions (Lai et al., 2024; Koh et al., 2024b; Pan et al., 2024). These agents are particularly effective in environments where human-like interaction with a user interface is necessary (Drouin et al., 2024b). Frameworks such as WebArena have further refined the evaluation of such agents by providing complex and realistic web navigation tasks (Zhou et al., 2023). Our work explores the Hybrid Agent that combine web browsing with API interactions. While prior work primarily focuses on browsing-only agents, we examine how Hybrid Agents can enhance performance by integrating structured API calls with web navigation.
Code Generation Agents and Tool Usage
Another stream of research focuses on agents that interact with online content via application programming interfaces (APIs) (Yuan et al., 2024; Patil et al., 2023; Wang et al., 2024d; b). In this context, works such as CodeAct have pioneered the development of agents that generate and execute code, including API calls, to perform tasks typically reserved for software engineers (Wang et al., 2024a; Zhang et al., 2024; Tang et al., 2024). These API-based agents are optimized for tasks that involve structured data exchanges, allowing them to perform operations more efficiently than traditional web navigation agents (Shen et al., 2024). On the other hand, our work integrates both browsing and API interactions, demonstrating that Hybrid Agents can outperform API-only agents in tasks requiring web navigation. While existing research shows the efficiency of API-based agents, our Hybrid Agent dynamically switches between APIs and web browsing to optimize task performance.
@@ 8 Conclusion and Future Work
In this paper, we propose new web agents that use APIs instead of traditionally browsers. We found that API-based agents outperform browsing-based counterparts, especially on websites with sufficient API support. Hence we further propose an agent that is capable of switching between using APIs or browsers and empirically outperforms agents that only uses one of the two interfaces.
For future work, we aim to explore methods for automatically inducing APIs using techniques such as Agent Workflow Memory (AWM) (Wang et al., 2024e). These methods could identify and generate API calls for websites lacking formal API support, further expanding the applicability and efficiency of API-based approaches. By automating the discovery and utilization of APIs, we envision even more robust agents capable of handling diverse web tasks with minimal reliance on manual interaction through browsing.
@@ 9 Limitations
In this paper, we evaluate web agents exclusively on WebArena tasks. While WebArena offers realistic and diverse challenges, the number and variety of tasks may be limited. Other benchmarks, such as MiniWoB (Shi et al., 2017), Mind2Web (Deng et al., 2023), and VisualWebArena (Koh et al., 2024a), provide alternative evaluation platforms. However, as discussed in Section 2.1, WebArena aligns more closely with real-world scenarios, while VisualWebArena is less applicable to our study because WebArena APIs lack support for interacting with images, a core component of VisualWebArena tasks.
@@
10 Acknowledgement
This work was supported in part by a grant from DSTA Singapore. The authors would like to thank CMU NeuLab colleagues for their constructive comments.

### v2

@@ 5.2 API Statistics for WebArena Sites
The API support for WebArena websites can be categorized into three levels: good, medium, and poor. APIs’ availability, functionality, and documentation, as described in Table 1, play a crucial role in the efficiency and flexibility of our agents101010See Appendix A.3 for where to find the WebArena APIs..
| Sites | Gitlab | Map | Shop | Admin | |
| # APIs | 988 | 53 | 556 | 556 | 31 |
| Quality | Good | Good | Fair | Fair | Poor |
@@ 5.2.1 Good API Support
Gitlab supports 988 endpoints, which offer extensive coverage across a wide range of functionalities, including repositories, commits, and users. This comprehensive API support allows for effective interaction in most WebArena tasks, making Gitlab one of the best-supported platforms in terms of API availability.
Map The Map website offers 53 endpoints. Despite the smaller number of endpoints, the APIs available are well-documented and cover most of the essential WebArena use cases.
@@ 5.2.2 Medium API Support
The Shopping and Shopping Admin websites share a common set of 556 APIs, which provide a reasonable level of support for common shopping tasks. However, some features, such as adding items to wish lists, are absent, and thus these tasks must be handled via browsing. Despite this, the documentation is fairly detailed. Overall, API calling is a solid, though not exhaustive, solution for handling shopping tasks.
@@ 5.2.3 Poor API Support
The WebArena Reddit is a self-hosted limited clone of the actual Reddit111111See Appendix A.3 for more explanations. with only 31 endpoints. It offers minimal API support and no documentation, making it the least API-friendly site in WebArena. Many critical functionalities such as searching posts are missing, significantly hampering task execution on Reddit, highlighting the need for a hybrid browsing+API approach.
@@ 5.3 API Implementation Details
We follow the methodologies discussed in Section 3.3 to provide APIs to agents. Appendix A.3 contains the sources of the public API documentations.
@@ 5.3.1 One-Stage Documentation for Small API Sets
For websites with fewer than 100 API endpoints, namely the Map and Reddit websites, we directly provide the full documentation to the agent.
Map The README documentation was inputted directly from the public API documentation.
Reddit Since there was no pre-existing documentation for the APIs, we leveraged GPT-4o121212https://openai.com/index/hello-gpt-4o/ itself to generate these README files. By prompting GPT-4o with a file containing all implementations of the API endpoints, we generated a README documentation, including input parameters, expected outputs, and example API calls.
@@ 5.3.2 Two-Stage Documentation Retrieval for Large API Sets
For websites with more than 100 endpoints, namely GitLab, Shopping, and Shopping Admin, we employ a two-stage documentation retrieval process.
We obtained Gitlab README documentations from the official website. For Shopping and Shopping Admin, the documentation is provided as OpenAPI specification, structured in YAML format.
@@ 5.4 Evaluation Framework
| WebArena Base (Zhou et al., 2024) | 15.0 | 15.6 | 13.9 | 10.4 | 6.6 | 8.3 | 12.3 |
@@
| Hybrid Agent | 44.4 | 45.9 | 25.7 | 41.2 | 51.9 | 16.7 | 38.9 |
We employed OpenHands as our evaluation framework to facilitate the development and testing of our agents (Wang et al., 2024c). OpenHands is an open-source platform designed for creating and evaluating AI agents that interact with both software and web environments, making it an appropriate infrastructure for our proposed methods. The OpenHands architecture supports various interfaces for agents to interact with. Moreover, this framework allows agents to keep a detailed record of past actions in the prompt, enabling agents to execute actions in a way that is consistent with earlier steps. For coding tasks, it implements an agent based on CodeAct (Wang et al., 2024a) that incorporates a sandboxed bash operating system and Jupyter IPython131313https://ipython.org environments, enabling Python code execution. Additionally, it includes a BrowsingAgent Browsing Agent that focuses solely on web navigation. This agent operates within a Chromium web browser powered by Playwright141414https://playwright.dev/, utilizing a comprehensive set of browser actions defined by BrowserGym (Drouin et al., 2024b). However, while the Browsing Agent can browse websites, and the CodeActAgent make API calls and execute code, there is not an agent that can natively do both. Given this base, we developed two varieties of agents for API-based solving of web tasks.
API-Based Agent
Our API-Based Agent essentially uses the CodeAct architecture (Wang et al., 2024a). In addition to the basic CodeAct framework, we tailor the agent for API calling by adding specialized instructions and examples that guide its understanding and using of APIs. At each step, the agent could utilize all previous actions to make informed selection of actions. The prompt of the API-Based Agent is included in the Appendix A.6.
Hybrid Browsing/API Calling Agent
In addition to the API-Based Agent, we developed a Hybrid Agent that integrates Chromium web browsing functionalities powered by Playwright into the existing API-Based Agent framework. This Hybrid Agent is provided the prompt describing both the APIs and the browsing actions, allowing for free transitions between API calling and web browsing. At each step, the agent can utilize the current state of the browser, all previous actions taken by the agent, and the results of those actions to determine the next course of action. The prompt of the Hybrid Agent is included in the Appendix A.7.
6 Results
6.1 Main Results
@@
The main results of our evaluation, as summarized in Table 2, demonstrate the performance of three different agents across WebArena websites.
@@
The API-Based Agent consistently achieved higher scores on most websites compared to the Browsing Agent. This agent’s strong performance is attributed to its specialized design for API calling, enabling it to efficiently interact with websites and complete tasks with no reliance on browsing.
@@
In contrast, the Browsing Agent, designed solely for navigating web interfaces, demonstrated significantly lower performance across most domains. It achieved its best scores on Shopping Admin and Map, but struggled more on the other websites.
@@
The Hybrid Agent, integrating both API calling and web browsing, outperformed the other agents on many websites. The agent’s ability to interleave API calling and web browsing proved beneficial. API calling delivers high performance for web tasks when well-supported APIs are available, while web browsing serves as a backup when API endpoints are unavailable or incomplete. Even if the website provides comprehensive APIs, there might be corner cases where APIs are not supportive. Thus, relying on web browsing is still needed for tasks that would otherwise fail through API-only interactions. Table 3 documents the frequency of each type actions of the Hybrid Agent: it chooses to do both Browsing and API in 77.7% of WebArena tasks, and it shows higher accuracy when choosing API only and API+browsing. More detailed analysis on action types, steps and cost and case studies are in Appendix A.4 and A.5.
| Actions | Frequency (%) | Accuracy (%) |
| Browsing only | 14.3 | 21.6 |
| API only | 8.0 | 38.5 |
| Browsing+API | 77.7 | 38.2 |
@@ 6.2 Does API Quality Matter?
Yes, API quality does significantly impact the performance of agents. High quality APIs provide comprehensive and well-documented endpoints that enable agents to interact accurately and efficiently with websites. With comprehensive API support, the API-Based Agent is able to tackle more tasks through API calling, while the Hybrid Agent rely less on browsing; on the other hand, clear and detailed documentation allows agents to use APIs effectively, ensuring that requests are accurate, and minimizing potential errors in task execution. For example, Gitlab and Map, with the best API support as mentioned in Section 5.2, demonstrate highest task completion accuracies among websites by the API-Based and Hybrid Agent.
Conversely, low-quality APIs, characterized by incomplete functionality or ambiguous documentation, can significantly degrade performance. In such cases, the absence of necessary endpoints may prevent the API-Based Agent from completing tasks and force the Hybrid Agent to resort to browsing. Moreover, poorly documented APIs can result in misusing parameters and headers, further reducing the effectiveness of the agent. This highlights the importance for websites to maintain comprehensive and well-documented API support.
@@ 7 Conclusion and Future Work
In this paper, we propose new web agents that use APIs instead of traditional browsers. We find that API-Based Agents outperform Browsing Agents, especially on sites with good API support. Thus we further propose an agent capable of interleaving API calling and browsing that empirically outperforms agents that only use one of the two interfaces.
For future work, we aim to explore automatically inducing APIs using methods such as Agent Workflow Memory (AWM) (Wang et al., 2024e). These methods could identify and generate API calls for websites lacking formal API support, further expanding the applicability and efficiency of API-Based Agents. By automating the discovery and utilization of APIs, we envision even more robust agents capable of handling diverse web tasks without reliance on interaction through browsing.
@@ 8 Limitations
In this paper, we evaluate web agents exclusively on WebArena tasks. While WebArena offers realistic and diverse challenges, the number and variety of tasks may be limited. Other benchmarks, such as Webshop (Yao et al., 2022), MiniWoB (Shi et al., 2017), Mind2Web (Deng et al., 2023), WebVoyager (He et al., 2024b), and VisualWebArena (Koh et al., 2024a), provide alternative valuable evaluation platforms. However, as discussed in Section 2.1, WebArena aligns more closely with real-world scenarios and our use case, while other benchmarks lack support for API calling. For example, VisualWebArena is less applicable to our study because WebArena APIs lack support for interacting with images, a core component of VisualWebArena tasks.